OpenStudy (anonymous):

Evaluate the expression (5x^2)^3/(10y)^4 for x=10 and y=5 A.15/128 B.60 C.5/256 D.20 Can someone show me how to do this please

OpenStudy (anonymous):

(5(10)^2)^3/(10(5))^4 =Just plug in the corresponding x and y values Then you have to do x squared before you multiply it by 5 (5(100))^3/(10(5))^4 Now solve for what is in the top parenthesis. (500)^3/(10(5))^4 500^3=125000000 Now work on the denominator. Do what is inside the parenthesis first 10*5=50 Now (50)^4=6250000 Now divide the numerator and the denominator. 125000000/6250000 Which equals 20 the answer is D

OpenStudy (lgbasallote):

substitute x = 10 and y = 5 \[\frac{(5x^2)^3}{(10y)^4}\] \[\frac{(5 \times 10^2)^3}{(10 \times 5)^4}\] \[\frac{(5 \times 100)^3}{(50)^4}\] \[\frac{500^3}{50^4}\] that is a big number

OpenStudy (anonymous):

thanks guys

OpenStudy (arabpride):

2 years later ~woosh~

OpenStudy (anonymous):

hahaha,arabpride you crack me up bro